\begin{thebibliography}{12}

\bibitem{GST}
M.Green, J.Schwarz and E.Witten, {\it Superstring Theory},
Cambridge University Press, 1987; \\
A.Polyakov, {\it Gauge Fields and Strings}, 1987; \\
J.Polchinsky, {\it String Theory}, Cambridge University Press, 1998;\\
A.Morozov, {\it String Theory: What is It?}
Sov.Phys.Usp. {\bf 35} (1992) 671-714
(Usp.Fiz.Nauk {\bf 162} 83 - 176)
\bibitem{Mth} P.K. Townsend, {\it The eleven-dimensional
supermembrane revisited}, Phys.Lett. {\bf B350} (1995) 184-187,
hep-th/9501068;\\
E.Witten, {\it String Theory Dynamics In Various Dimensions},
Nucl.Phys. {\bf B443} (1995) 85-126, hep-th/9503124
\bibitem{BL} J.Bagger and N.Lambert, {\it Modelling Multiple M2's},
hep-th/0611108 v3;
{\it Gauge Symmetry and Supersymmetry of Multiple M2-Branes},
arXiv: 0711.0955;
{\it Comments on Multiple M2-Branes}, arXiv: 0712.3738
\bibitem{G} A.Gustavsson, {\it Algebraic Structures
on Parallel M2-branes}, arXiv: 0709.1260 v5
\bibitem{JS1} J.Schwarz, {\it Superconformal
Chern-Simons Theories}, JHEP {\bf 0411} (2004) 078,
hep-th/0411077
\bibitem{Po} A.Polyakov, {\it Quantum Geometry of Bosonic String},
Phys.Lett. {\bf 103B} (1981) 207-210;
{\it Quantum Geometry of Fermionic String},
Phys.Lett. {\bf 103B} (1981) 211-213
\bibitem{BK} A.Belavin and V.Knizhnik,
{\it Algebraic Geometry and the Geometry of Quantum Strings},
Phys.Lett. {\bf 168B} (1986) 201-206;  ZhETF {\bf 91} (1986) 247;\\
A.Belavin, V.Knizhnik, A.Morozov and A.Perelomov,
{\it Two- and three-loop amplitudes in the bosonic string theory},
Phys.Lett. {\bf 177B} (1986) 324; Pis'ma v ZhETF,
{\bf 43} (1986) 319;\\
G.Moore, {\it Modular Forms and Two-Loop String
Physics}, Phys.Lett. {\bf 176B} (1986) 69; \\
A.Beilinson and Yu.Manin, {\it The Mumford form and the Polyakov
measure in string theory},
Comm. Math. Phys. {\bf 107} (1986) 359-376;\\
A.Morozov, {\it Explicit formulae for one, two, three and four
loop string amplitudes}, Phys.Lett. {\bf 184B} (1987) 171-176;
{\it Analytical Anomaly and Heterotic String in the Formalism
of Continual Integration}, Phys.Lett. {\bf 184B} (1987) 177-183
\bibitem{BKoth}
L.Alvarez-Gaume, J.Bost, G.Moore, P.Nelson and
C.Vafa, {\it Bosonisation on Higher Genus Riemann Surfaces},
Comm.Math.Phys. {\bf 112} (1987) 503;
{\it Modular forms and the cosmological constant},
Phys.Lett. {\bf 178B} (1986) 41; \\
E.Verlinde and H.Verlinde,
{\it Chiral Bosonization, Determinants and the String Partition
Function}, Nucl. Phys. {\bf B288} (1987) 357-396; \\
E.D'Hoker and D.Phong, Nucl.Phys. {\bf B292} (1987) 109;
{\it The Geometry of String Perturbation Theory}, Rev.Mod.Phys.,
{\bf 60} (1988) 917-1065; \\
{\it Multiloop Amplitudes in the Theory of Quantum Strings
and Complex Geometry}, Usp. Phys. Nauk, {\bf 159} (1989) 401-453; \\
A.Gerasimov, A.Marshakov, A.Morozov, M.Olshanetsky and S.Shatashvili,
{\it Wess-Zumino-Witten model as a theory of free fields},
Int.J.Mod.Phys. {\bf A5} (1990) 2495-2589
\bibitem{BLGfirst} A.Gustavsson, {\it Selfdual Strings and
Loop Space Nahm Equations}, arXiv: 0802.3456 v2
\bibitem{Mu} S.Mukhi and C.Papageorgakis,
{\it M2 to D2}, arXiv: 0803.3218
\bibitem{BLS} M.Bandres, A.Lipstein and J.Schwarz,
{\it N=8 Superconformal Chern-Simons Thepries},
arXiv: 0803.3242
\bibitem{BTT} D.Berman, L.Tadrowski and D.Thompson,
{\it Aspects of Multiple Branes}, arXiv: 0803.3611
\bibitem{MVR} M. Van Raamsdonk,
{\it Comments on the Bagger-Lambert Theory and
Multiple M2-Branes}, arXiv: 0803.3803
\bibitem{M} A.Morozov, {\it On the Problem of Multiple M2 Branes},
arXiv: 0804.0913
\bibitem{Gu} A.Gustavsson, {\it private communication}
\bibitem{DMP} J.Distler, S.Mukhi, C.Papageorgakis and
M.Van Raamsdonk, {\it M2-branes on M-folds}, arXiv:0804.1256
\bibitem{Gra} U.Gran, B.E.W. Nilsson and C.Petersson,
{\it On relating multiple M2 and D2-branes},
arXiv:0804.1784
\bibitem{HH} P.-M.Ho, R.-C.Hou and Y.Matsuo,
{\it Lie 3-Algebra and Multiple M2-branes}, arXiv:0804.2110
\bibitem{GSP} J.Gomis, A.Jafari Salim and F.Passerini,
{\it Matrix Theory of Type IIB Plane Wave from Membranes},
arXiv:0804.2186
\bibitem{BRH} E.A.Bergshoeff, M.de Roo and O.Hohm,
{\it Multiple M2-branes and the Embedding Tensor}, arXiv:0804.2201
\bibitem{HLL} K.Hosomichi, K.-M.Lee and S.Lee,
{\it Mass-Deformed Bagger-Lambert Theory and its BPS Objects},
 arXiv:0804.2519
\bibitem{Pa} G. Papadopoulos,
{\it M2-branes, 3-Lie Algebras and Plucker relations},
arXiv:0804.2662;
{\it On the structure of k-Lie algebras}, arXiv:0804.3567
\bibitem{GG} J.Gauntlett and J.Gutowski,
{\it Constraining Maximally Supersymmetric Membrane Actions},
arXiv:0804.3078
\bibitem{BLGlast}
P.-M.Ho and Y.Matsuo, {\it M5 from M2}, arXiv:0804.3629
\bibitem{triple}
J.Gomis, G.Milanesi, and J.G.Russo,
{\it Bagger-Lambert Theory for General Lie Algebras},
arXiv:0805.1012 v2; \\
S.Benvenuti, D.Rodriguez-Gomez, E.Tonni and H.Verlinde,
{\it N=8 superconformal gauge theories and M2 branes},
arXiv:0805.1087;\\
P.-M.Ho, Y.Imamura and Y.Matsuo,
{\it M2 to D2 revisited}, arXiv:0805.1202
\bibitem{3alg}
V.Filipov, Sib.Math.Jour. {\bf 26} No.6
(1985) 126; \\
L.Takhtajan, Comm.Math.Phys. {\bf 160} (1994) 295,
hep-th/9301111
\bibitem{triNa} H.Awata, M.Li, D.Minic and T.Yoneya,
{\it On the Quantization of Nambu Brackets},
JHEP {\bf 0102} (2001) 013, hep-th/9906248
\bibitem{Kro} D.Krotov and A.Morozov,
{\it A Solvable sector of AdS theory}, JHEP 0510:062,2005,
hep-th/0506028
\bibitem{LM} A.Levin and A.Morozov, {\it On the Foundations
of the Random Lattice Approach to Quantum Gravity}, Phys. Lett.
{\bf 243B} (1990) 207-214
\bibitem{BDH} L.Brink, P.Di Vecchia and P.Howe,
Phys.Lett. {\bf B65} (1976) 471
\bibitem{mamos}
F.David, {\it Planar diagrams, two-dimensional lattice gravity
and surface models}, Nucl.Phys. {\bf B257} (1985) 45;
{\it A model of random surfaces with non-trivial critical behavior},
Nucl.Phys. {\bf B257} (1985) 543; \\
J.Ambjorn, B.Durhuus and J.Frohlich, {\it Deseases of triangulated
random surface models and possible cures}, Nucl.Phys. {\bf B257}
(1985) 433; \\
V.Kazakov, I.Kostov and A.Migdal, {\it Critical properties
of randomly triangulated planar random surfaces},
Phys.Lett. {\bf B157} (1985) 295-300; \\
D.-J.Smit, {\it Summations over equilatteraly triangulated
surfaces and the critical string measure}, Comm.Math.Phys.
{\bf 143} (1992) 253-286; \\
A.Morozov, {\it Integrability and Matrix Models}, Phys.Usp. {\bf 37}
(1994) 1-55, hep-th/9303139; hep-th/9502091
\bibitem{Gro} G.Shabat and V.Voevodsky, {\it Drawing curves
over number fields}, in: "The Grothendieck Festschrift, v.III",
Prog.in Math., Birkhauser, 1990, pp.199-227
\bibitem{Lisi} A.G.Lisi,
{\it Clifford bundle formulation of BF gravity generalized
to the standard model}, gr-qc/0511120;
{\it An Exeptionally Simple
Theory of Everything}, arXiv: 0711.0770
\bibitem{GSA}
M.Green and J.Schwarz,
{\it Covariant Description of Supesrstrings}, Phys.Lett.
{\bf 136B} (1984) 367-370; \\
S.Carlip, Nucl.Phys. {\bf B284} (1987) 365;
Phys.Lett. {\bf 186B} (1987) 141;\\
R.Kallosh and A.Morozov, {\it Green-Schwarz action
and loop calculations for superstring}, Int.J.Mod.Phys. {\bf A3}
(1988) 1943-1958;
{\it On the vanishing of multiloop contributions to 0,1,2,3-point
functions in the Green-Schwarz formalism for heterotic string},
Phys.Lett. {\bf 207B} (1988) 164-168;\\
N.Berkovits, {\it Explaining the Pure Spinor
Formalism for the Superstring}, arXiv: 0712.0324
\bibitem{DK}
J.Polchinsky, {\it A two-dimensional model for quantum gravity},
Nucl.Phys. {\bf B234} (1989) 123; \\
F.David, {\it Conformal field theories coupled to 2d gravity
in the conformal gauge}, Mod.Phys.Lett. {\bf A3} (1988) 1651; \\
J.Distler and H.Kawai, {\it Conformal field theory and 2d gravity or
who's afraid of Joseph Liouville}, Nucl.Phys. {\bf B312} (1989) 509
\bibitem{Ie}
R.Iengo and C.J.Zhu,
{\it Explicit modular invariant two-loop super-string
amplitude relevant for $R^4$}, JHEP {\bf 06} (1999) 011; \\
R.Iengo, {\it Computing the $R^4$ term at two super-string loops},
JHEP {\bf 02} (2002) 035
\bibitem{DHP} E.D'Hoker and D.Phong, {\it Two Loop Superstrings},
Phys.Lett. {\bf B529} (2002) 241-255, hep-th/0110247;
Nucl.Phys. {\bf B636} (2002) 3-60, hep-th/0110283;
Nucl.Phys. {\bf B636} (2002) 61-79, hep-th/0111016;
Nucl.Phys. {\bf B639} (2002) 129-181, hep-th/0111040;
Nucl.Phys. {\bf B715} (2005), 3-90, hep-th/0501197;
Nucl.Phys. {\bf B715} (2005), 91-119, hep-th/0501196;\\
S.Cacciatori, and F.Dalla Piazza,
{\it Two loop superstring ampliudes and $S_6$ representations},
arXiv: 0707.0646; \\
S.Cacciatori, F.Dalla Piazza and B.van Geemen,
{\it Modular Forms and Three Loop Superstring Amplitudes},
arXiv: 0801.2543;
{\it Genus Four Superstring Measures}, arXiv: 0804.0457 v1;\\
F.Dalla Piazza and B.van Geemen,
{\it Siegel Modular Forms and Finite Symplectic Groups},
arXiv: 0804.3769;
{\it Lectures on
Two-Loop Superstrings}, hep-th/0211111; \\
S.Grushevsky, {\it Superstring Amplitudes
in Higher Genus}, arXiv: 0803.3469; \\
R.Salvati Manni, {\it Remarks on Superstring
Amplitudes in Higher Genus}, arXiv: 0804.0512; \\
A.Morozov, {\it NSR Superstring Measures Revisited},
arXiv: 0804.3167 v2;
{\it NSR measures on hyperelliptic locus and
non-renormalization of 1,2,3-point functions}, arXiv: 0805.0011
\bibitem{rtheta}
J.Fay, {\it Theta functions on Riemann surfaces},
Lect.Notes Math. {\bf 352}, Springer, 1973; \\
D.Mumford, {\it Tata Lectures on Theta}, Progr.in Math.
{\bf 28, 43}, Birkhauser, 1983, 1984
\bibitem{Ok} T.Hollowood, A.Iqbal and C.Vafa,
{\it Matrix models, geometric engineering and elliptic genera},
hep-th/0310272; \\
A.Okounkov, {\it Random partitions and instanton counting},
math-ph/0601062







\end{thebibliography}

\documentstyle{article}

\def\be{\begin{eqnarray}}
\def\ee{\end{eqnarray}}
\def\nn{\nonumber}

\hoffset= - 1.0in         

\begin{document}

\hfill ITEP/TH-25/08

\bigskip

\centerline{\Large{
From Simplified BLG Action to the First-Quantized M-Theory
}}

\bigskip

\centerline{A.Morozov}

\bigskip

\centerline{\it ITEP, Moscow, Russia}

\bigskip

\centerline{ABSTRACT}

\bigskip

{\footnotesize
Concise summary of the recent progress in the search for
the world-volume action for multiple M2 branes.
After a recent discovery of simplified version of BLG
action, which is based on the ordinary Lie-algebra
structure, does not have coupling constants and
extra dynamical fields, attention should be switched
to the study of M2 brane dynamics.
A viable brane analogue of Polyakov formalism and
Belavin-Knizhnik theorem for strings can probably
be provided by Palatini formalism for 3d (super)gravity.
}

\bigskip

\bigskip

In the context of general string theory \cite{GST}
a variety of string and superstring models,
linked by a number of duality relations, is naturally
unified in a hypothetical "M-theory" \cite{Mth},
which in its most naive perturbative phase is represented
by "fundamental (super)membranes".
The BLG action \cite{BL,G} resolves the long-standing
puzzle \cite{JS1} of finding the 3d Lagrangian with
appropriate superconformal symmetry and thus opens a way
to developing the first-quantized theory of M2 branes.
This implies that M-theory can now be studied in the same
constructive manner as bosonic and super-strings in 1980's
\cite{Po,BK,BKoth}.
Naturally, this ground-breaking achievement attracts enormous
attention \cite{BLGfirst}-\cite{triple}, and some minor drawbacks
of original analysis in \cite{BL,G} are now fully cured.

The main obvious difficulty of original BLG formulation
was its reliance upon sophisticated 3-algebra
(quantum Nambu bracket) structure \cite{3alg}
-- new for the fundamental physical considerations.
The lack of experience and intuition about this structure
caused certain confusion at the early stages:
original BLG action was written only for an artificial(?)
example of $SO(4)$ symmetry, problems were discovered with
straightforward generalizations to other groups and even
doubts appeared about the very existence of BLG action for
the stack of $N$ M2 branes with arbitrary $N$, which would
require promotion of $SO(4)$ to $SU(N)$.
The key step in overcoming this problem was analysis of the
M2 $\rightarrow$ D2 conversion in \cite{Mu}, which linked
the 3-algebra structure to conventional Lie algebras,
governing Yang-Mills and D-brane theories.
Based on this analysis, in \cite{M} a "simplified" BLG action
was introduced, which makes use of the Lie-algebra structure
only (i.e. is based on "reducible-to-Lie-algebra" Nambu bracket
of \cite{triNa}, see eq.(\ref{Nabr}) below).
The only new ingredient, distinguishing this version of
BLG action from the ones familiar from string/brane studies
was a pair of extra color-less octuplets $\varphi^I$
and $\chi^A$.
While very simple, this suggestion had serious problems as it was,
originated from degeneracy of the underlying Nambu bracket
and the lack of total antisymmetry of 3-algebra structure constants:
this made original supersymmetry proof of \cite{BL} unapplicable
and the action in \cite{M} potentially non-supersymmetric.
Thus it was meant to be a toy-example, showing the direction
to eliminate unnecessary(?) elements of the BLG construction,
but still possessing some extra fields and requiring some further
tuning.
A natural next step was to look at a central extension,
lifting degeneracy of Nambu bracket \cite{Gu}
-- and this was finally done in a triple of wonderful papers
\cite{triple}.
They resolved the discrepancy between \cite{M} and BLG approach
in an  elegant way, by changing the nature of the extra
fields $\varphi,\chi$: they are actually auxiliary, non-dynamical
variables.
Kinetic terms
$(\partial_\mu\varphi^I)^2$ and $\bar\chi^A\hat\partial\chi^A$
of \cite{M} are substituted in \cite{triple} by
$\partial_\mu\tilde\varphi^I\partial_\mu\varphi^I$ and
$\frac{1}{2}\bar{\tilde\chi}^A\hat\partial\chi^A
+ \frac{1}{2}\bar{\chi}^A\hat\partial\tilde\chi^A$
where $\tilde\varphi,\tilde\chi$ is still another
pair of color-less octuplets which do not appear anywhere else
in the action and thus serve as Lagrange multipliers,
eliminating the fluctuations of the unwanted
$\varphi$ and $\chi$ fields.
In other words, the modified version of the simplified BLG
action of \cite{M} is now \cite{triple}:
\be
-\frac{1}{2}{\rm tr}\Big(D_\mu \phi^I - B_\mu\varphi^I\Big)^2
+\frac{i}{2} {\rm tr} \psi^A\hat D \Big(\psi^A - \hat B\chi^A\Big)
+ \nn \\ +
\Big(\partial_\mu\tilde\varphi^I - {\rm tr}( B_\mu \phi^I)\Big)
\partial_\mu\varphi^I
 -\frac{i}{2}\bar{\tilde\chi}^A \hat\partial\chi^A -
\frac{i}{2}\bar\chi^A\Big(\hat \partial\tilde\chi^A -
{\rm tr}(B_\mu \psi^A)\Big) + \nn \\
+ \frac{1}{2}\epsilon^{\mu\nu\lambda} {\rm tr}
\Big(F_{\mu\nu}B_\lambda\Big) -
\frac{1}{12}{\rm tr} \Big(\varphi^I[\phi^J,\phi^K] +
\varphi^j[\phi^K,\phi^I] + \varphi^K[\phi^I,\phi^J]\Big)^2 +\nn\\
+ \frac{i}{2}\Gamma_{AB}^{IJ}\varphi^I
{\rm tr}\Big(\bar\psi^A[\phi^J,\psi^B]\Big)
+\frac{i}{4}\Gamma^{IJ}_{AB} {\rm tr}\Big(\bar\psi^A
[\phi^I,\phi^J]\Big)\chi^B
-\frac{i}{4}\Gamma_{AB}^{IJ} \bar\chi^A {\rm tr}
\Big([\phi^I,\phi^J]\psi^B\Big)
\label{siBLG}
\ee
It essentially differs from eq.(21) of \cite{M} in the second
line.\footnote{A trivial mistake of \cite{M}
in the $\phi^6$ term (omitted item $2\sum_{I<J}\varphi^I\varphi^J
{\rm tr}\Big([\phi^I,\phi^K][\phi_j,\phi^K]\Big)$
is also corrected in \cite{triple} and in (\ref{siBLG}).}
Here $\phi^I$ and $\psi^A$ with $I=1,\ldots,8$ and $A=1,\ldots,8$
are real- and Grassmann-valued elements of the vector and spinor
representations of the $SO(8)$ group respectively (related by
octonionic triality to the second spinor representation, where
the $N=8$ SUSY transformation parameter takes values).
They are also $N\times N$ matrices, i.e. belong to adjoint
representation of the gauge group $G=SU(N)$.
$A_\mu$ is the corresponding connection, also in the adjoint of $G$,
$D_\mu \phi = \partial_\mu\phi - [A_\mu,\phi]$, $F_{\mu\nu} =
\partial_\mu A_\nu - \partial_\nu A_\mu + [A_\mu,A_\nu]$,
and $B_\mu$ is an auxiliary adjoint vector field (not a connection).
The other auxiliary fields $\varphi^I,\tilde\varphi^I$ and their
superpartners $\chi^A,\tilde\chi^A$ are $G$-singlets (are color-less),
possibly fragments of Kac-Moody extension of $G$.
See \cite{M,triple} for further details.

\bigskip

As explained in \cite{triple},

$\bullet$ The action (\ref{siBLG})
is $N=8$ supersymmetric due to original BL theorem \cite{BL},
because it is now based on the 3-algebra with totally
antisymmetric structure constants, which is a central extension
of the degenerate one \cite{triNa} used in \cite{M}:
\be
\left[X,Y,Z\right] = {\rm tr}\, X \cdot [Y,Z] +
{\rm tr}\, Y \cdot [Z,X] + {\rm tr}\, Z \cdot [X,Y] +
\zeta \cdot {\rm tr}\left( X [Y,Z]\right)
\label{Nabr}
\ee
$\zeta$ is a central element, different from unity matrix $I$
and related to it by non-trivial scalar product
$<\!I,I\!>\, = \,<\!\zeta,\zeta\!>\, =0$,
$\ <\!I,\zeta\!>\, =\, <\!\zeta,I\!>\, = -1$,
so that the 3-algebra metric is
$\left(\begin{array}{cc|c}
0&-1&0\\-1&0&0\\ \hline  0&0& h
\end{array}\right)$ and 
$f^{abcI} = -{f^{abc}}_\zeta = -f^{abc} = -f^{Iabc}$.
The last term of (\ref{Nabr}) was absent in \cite{M}
and this made the 3-bracket degenerate and the structure constants
(with the forth index raised by 3-algebra metric)
not totally antisymmetric.
The $\varphi,\chi$ fields are associated with the $I$
(matrix-trace) generator, while $\tilde\varphi,\tilde\chi$
-- with the central element $\zeta$.
Non-trivial 3-algebra metric implies the $\varphi$-$\tilde\varphi$
and $\chi$-$\tilde\chi$ mixing form of the kinetic terms in
(\ref{siBLG}).

$\bullet$ The would-be coupling constant in front of
non-quadratic terms in (\ref{siBLG}) can be absorbed into
rescaling of $\varphi$ and $\chi$, accompanied by rescaling
of $\tilde\varphi$ and $\tilde\chi$ in the opposite direction.
This lack of this feature was one of the problems in \cite{M},
and in (\ref{siBLG}) we have an action, which has no dimensionless
coupling constants, as required in M-theory.

$\bullet$ All the unwanted extra fields
$\varphi,\chi,\tilde\varphi,\tilde\chi$ are auxiliary:
they do not propagate and contribute only through boundary
terms\footnote{In this respect the action in the
$\varphi$-$\tilde\varphi$ sector is reminiscent of the one,
considered in \cite{Kro}.}
(i.e. to correlators) and zero-modes.

$\bullet$ The fact that Lagrange multiplier $\tilde\varphi$
nullifies only $\partial^2\varphi$ rather than $\varphi$ itself
is very important, because this allows the zero-mode
$\varphi=const$.
Among other effects, this zero mode can form
a condensate, producing a term
$<\!\varphi\!>^2\!{\rm tr}\, B^2$ 
from the first item in (\ref{siBLG}),
which, after auxiliary field $B_\mu$ is integrated away,
converts the Chern-Simons interaction ${\rm tr} F\wedge B$
into kinetic Yang-Mills term
$<\!\varphi\!>^{-2}\! {\rm tr}\, F_{\mu\nu}^2$ for connection $A_\mu$.
This means that despite $\varphi$ fields are now auxiliary,
the crucially important possibility to use them for the
M2 $\rightarrow$ D2 conversion {\it a la} \cite{Mu} is preserved.

\bigskip

All this means that today we possess a perfectly simple version
(\ref{siBLG}) of the BLG action for arbitrary number of M2
branes, there are no longer doubts about its existence for
arbitrary gauge group $SU(N)$, there are no coupling constants,
no extra dynamical fields, and it is clearly related to the other
brane actions, as required by embeddings of $d=10$ superstring
models into the $d=11$ M-theory.
The road is now open for building up the first-quantized theory
of M2 branes (supermembranes).
This implies that attention can now be shifted from the study of
3-algebra structure (where a lot of interesting questions still
remain) to the other issues:
we know what should be the crucial next steps from the history
of first-quantized theory of superstrings.

Constructing the action (\ref{siBLG}) can be considered 
as the very first step,
corresponding to substitution of Nambu-Goto action for bosonic
strings by a $\sigma$-model action, of which (\ref{siBLG})
is supposed to be a (super)membrane analogue.
In the case of membranes
the problem was more complicated, because Nambu-Goto action is
ill (does not damp fluctuations) from the very beginning,
no approach to bosonic membrane is still available (problems look
more severe than the tachyon of bosonic string) and one should
begin directly from the supersymmetric case, moreover supersymmetry
should be immediately extended to ${\cal N}=8$.
Thus it may be not too surprising that we had to wait till
2008 to have this action written down...

In the case of strings the next big step was consideration of
world sheets with non-trivial topologies, with two complementary
formalisms finally developed for this purpose
(and still not fully related,
see \cite{LM} for description of the corresponding problems).
One is the Polyakov formalism \cite{Po},
promoting the $\sigma$-model action to arbitrary curved
$2d$ geometries and generalizing the treatment of relativistic
particle in \cite{BDH}.
Another is equilateral-triangulation approach, nicely expressed
in terms of matrix models \cite{mamos}
and formally equivalent to substitution
of smooth $2d$-geometries by Grothendieck's
{\it dessins d'enfants} \cite{Gro}.
In the case of membranes this step is going to be a hard exercise,
already because the topological classification of $3d$ world
volumes is far more complicated than in $2d$.
Still, the very first movement -- introduction of $3d$ geometry 
into (\ref{siBLG}) by both above-mentioned methods --
should be straightforward, and undoubtedly very interesting.
For a variety of reasons it seems natural to do this in the
modern BF-version of Palatini formalism, which is now widely
popularized by controversial, but inspiring papers of G.Lisi
\cite{Lisi}.
Of certain help can be also comparison with the Green-Schwarz
formalism for the superstrings \cite{GSA}, where world-sheet
action has some common features with (\ref{siBLG}): it also
looks non-linear, but actually non-linearities concern only the
zero-modes and boundary effects.

Of crucial importance should be identification of the relevant
world-sheet-geometry degrees of freedom (moduli), which
the action is going to depend upon. This is not the $3d$
metric or dreibein and spin-connection themselves -- already
because of the general covariance. However, as we know
from experience with strings, the remaining degrees of freedom
(Liouville field) can also be irrelevant (or identified with
the other physical fields \cite{DK}), so that the only remaining
moduli are those of the $2d$ complex structures: finitely many
for any given $2d$ topology.
It is the analogue of Belavin-Knizhnik theorem \cite{BK,BKoth}
that formulates this statement for strings, which should be the
next big discovery in the story of BLG actions.
Again, there are many complications in the case of membranes:
as already mentioned, from the very beginning we need
supersymmetry (and the corresponding problem for superstrings was
partly resolved only quite recently! \cite{Ie,DHP}). Moreover,
the analogue of Riemann theta-function theory \cite{rtheta}
in $3d$ is not yet at our disposal -- and here we should face
the same problems as the other approaches to $3d$ topological
theories \cite{Ok}: there are no conventional terms to express
our answers through...

In this short summary we do not speculate about the resolution
of all these problems, i.e. about filling the empty spaces
in the right column of the following table:

\bigskip

\centerline{
\begin{tabular}{|c|c|}
\hline & \\
(super)strings & (super)membranes \\
&\\
\hline\hline & \\
$2d$ Nambu-Goto action $\rightarrow$ $2d$ $\sigma$-model action &
spirit of membrane $\rightarrow$
simplified BLG action (\ref{siBLG})\\
&\\
\hline & \\
Polyakov formalism:   & BF-version of Palatini formalism in $3d$\\
introduction of $2d$ metric, &\\
critical dimensions (where massless excitations occur), &\\
sum over geometries, & \\
sum over topologies,   & \\
relation to equilateral triangulations approach & \\
&\\
\hline & \\
Belavin-Knizhnik theorem: & \\
reduction of sum over metrics to sum over moduli &\\
&\\
\hline & \\
topology of world sheet:  &\\
spin structures and GSO projection, &\\
string field theory, &\\
boundary correlators and AdS/CFT correspondence & \\
&\\
\hline & \\
topology of the target space (compactifications): &\\
generic $2d$ conformal theories, &\\ $T$-dualities, &\\
other dualities & \\
&\\
\hline & \\
\ldots & \\
&\\
\hline
\end{tabular}
}

\bigskip \bigskip

\noindent
Our goal is to emphasize that we are
now in front of the new and interesting breakthrough into
the unknown -- the possibility opened to us by the timely
formulated problem \cite{JS1}, a brilliant insight \cite{BL,G}
and qualified polishing \cite{BLGfirst}-\cite{BLGlast},
culminated in \cite{triple} in the elegant formula 
(\ref{siBLG}), which is going to be -- perhaps, in some
reshaped and redecorated version -- a new focus of attention
in string theory in the coming years.

\section*{Acknowledgements}

This work  is partly supported by Russian Federal Nuclear
Energy Agency and Russian Academy of Sciences,
by the joint grant 06-01-92059-CE,  by NWO project 047.011.2004.026,
by INTAS grant 05-1000008-7865, by ANR-05-BLAN-0029-01 project,
by RFBR grant 07-02-00645 and
by the Russian President's Grant of Support for the Scientific
Schools NSh-3035.2008.2

\end{document}